Electro Company is planning to develop a new electric motorcycle. It estimates that it will need to invest $30 million in equipment and facilities. Its management team estimates that the new motorcycle will generate a net cash flow of $5 million in year 1, $7 million in year 2, $10 million in year 3, $14 million in year 4, and $17 million in year 5. Assuming that Electro evaluates all investments using a 15% rate of return (i.e., discount rate), what is the NPV of this project? Should it proceed? Show your work, and explain the reasoning for your answer.
Answer #1

А YEAR CASHFLOW (in million) -30 o aut WNW UWN PVF @ 15% Present Value -30 0.8695652174.347826087 0.756143667 5.293005671 0.6AB D FORMULA USED YEAR CASHFLOW (in million) PVF @ 15% | Present Value -30 =1/(1.15^B3 =C3*D3 =1/(1.15)^B4 =C4*D4 =1/(1.15)^B